INFRARED SPECTROSCOPY OF (CH_{3})_{3}N-H^{+}-(H_{2}O)_{n} (n = 1-22)

\underline{RYUNOSUKE SHISHIDO}, ASUKA FUJII, Department of Chemistry, Graduate School of Science, Tohoku University, Sendai 980-8578, Japan; and KUO JER-LAI, Institute of Atomic and Molecular Sciences Academia Sinica, Taipei, Taiwan.

The magic number behavior of H^{+}(H_{2}O)_{21} has been well known. Futhermore, this magic number is observed even in the mixed clusters H^{+}(H_{2}O)_{m}(NH_{3})_{n} and H^{+}(H_{2}O)_{m}(MeOH)_{n} (m + n = 21). This means that NH_{3} or MeOH molecules are compatible with water molecules in the hydrogen bond network of the magic number cluster. In the present study, infrared spectroscopy is applied to (CH_{3})_{3}N-H^{+}-(H_{2}O)_{n} (n = 1-22), and structures of these clusters are determined with help of density functional theory calculations. As a result, it is demonstrated that no magic number is seen in the case of (CH_{3})_{3}N-H^{+}-(H_{2}O)_{n}. (CH_{3})_{3}N is not a spectator to the hydrogen bond network of protonated water clusters, and it is largely changes the network structure.